OpenStudy (anonymous):

Logo function 3(5)2x−3^ = 6 Any assistance?

OpenStudy (anonymous):

\[3(5)^{2x-3} = 6\]

OpenStudy (anonymous):

First, divide both sides by 3. 5^(2x-2)=2. Take log 5 of both sides, and solve.

OpenStudy (anonymous):

Can you write that out?

OpenStudy (anonymous):

Sure. Here are the steps: \[5^{2x-2} = 2\] \[2x-2 = \log_5 2 =0.431\] \[2x = 2.431\] \[x = 1.2155 \]

OpenStudy (anonymous):

Wow ok Thank You, very helpful

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

Er wait...one problem, but solved :)

OpenStudy (anonymous):

was it suppose to be 2x-3 instead of 2x-2